OpenStudy (anonymous):

x2 + y2 = 25 y − x = 7

OpenStudy (luigi0210):

Solving?

OpenStudy (anonymous):

yeah!

OpenStudy (uri):

lol

OpenStudy (luigi0210):

solve for y in the second and plug it in.

OpenStudy (uri):

i knw the ansur.

OpenStudy (anonymous):

x+7 and then?

OpenStudy (luigi0210):

Yes, plug that in. @uri finish this for me. I gtg. And don't give the ansur xD

OpenStudy (johnweldon1993):

@wequess2 right...when you solve for 'y' in the 2nd...you get y = x + 7 Now you go back to your first equation \[\large x^2 + y^2 = 25\] And you now replace 'y' here with what you now know 'y' equals... \[\large x^2 + (x + 7)^2 = 25\] Now you simplify that....what is (x + 7)²?